OpenStudy (anonymous):

What two factors affect the magnitude of the gravitational force between two objects?

OpenStudy (abhisar):

Grvitational force between two objects is given by the formula \(\sf \boxed {\huge F_g=\huge \frac{G\times M_1\times M_2}{r^2}}\) Here, \(\square\) G= gravitational constant \(\square~\sf M_1~and~M_2\) are the masses of two objects. \(\square\) r is the distance between two objects. Now, G is a constant and only mass and distance between the two objects are variables which can effect the force of gravity. So what should be the answer to your question?

OpenStudy (abhisar):

Also,, I should have posted this question in Physics section :-)